A​ number, one fifth of that​ number, and one fourth of that number are added. The result is 58. What is the original​ number?
Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

A​ number, one fifth of that​ number, and one fourth of that number are added. The result is 58. What is the original​ number?
---------
n*(1 + 0.2 + 0.25) = 58

Answer by greenestamps(13334)



The sum is a whole number, so 1/5 of the number and 1/4 of the number are almost certain to be both whole numbers.

That means the number is divisible by both 4 and 5; and that means it is divisible by 4*5=20.

Then since the sum is 58, the number is probably the largest multiple of 20 less than 58.

So try 40 as the number:
40+40/5+40/4 = 40+8+10 = 58